OpenStudy (anonymous):

The concentration C of a certain drug in the bloodstream t hours after injection into muscle tissue is given by c (t) = (4t) / (2+t^3)

OpenStudy (anonymous):

when is concentration the greatest? thanks! :)

OpenStudy (abb0t):

Do you know derivatives?

OpenStudy (anonymous):

yes, i was able to get the first derivative but i got stuck at the second one f'(x)=(-8t^3 + 8)/(2+t^3)^2

OpenStudy (abb0t):

Why are you taking the second derivative?

OpenStudy (anonymous):

I thought that would be the best way to find the absolute max

OpenStudy (abb0t):

You can use your first derivative to find the max. Find the root(s). Then plug in the value for c(t) and that is your max.

OpenStudy (abb0t):

find when is at max concentration*

OpenStudy (anonymous):

by roots do you mean the result of setting f'(x)=0? I got the cube root of 16, so is that the only thing i need to plug in to the original equation?

OpenStudy (abb0t):

Yes, when f'(x) = 0

OpenStudy (abb0t):

How did you get \(\sqrt[3]{16}\)????

OpenStudy (abb0t):

\(-8t^3+8=0\) gave you \(t= \sqrt[3]{16}\)??

OpenStudy (anonymous):

oh, it should have been 1

OpenStudy (anonymous):

when i plugged in 1 to the original equation i got 4/3 but it wasnt the correct answer. . .

OpenStudy (tkhunny):

You have this: f'(x)=(-8t^3 + 8)/(2+t^3)^2 Why one earth are you playing with that -8? The denominator never is zero, so focus on the numerator. -8 is never zero, so get rid of it. t^3 - 1 = 0 And the only Real solution is apparent. Simplify your life. Don't make it unnecessarily complicated.

OpenStudy (tkhunny):

c(1) = 4/3, but why do we care? The question is WHEN, not HOW MUCH. We are done when we say t = 1 holds the greatest concentration.

OpenStudy (anonymous):

thank you for your help I appreciate it

OpenStudy (tkhunny):

Note: It is generally considered bad form to change variable names mid problem. If the concentration is C, it is not convenient to define the function c(t). A better choice would be C(t).

OpenStudy (tkhunny):

You should also check to see if it really is a max or min or neither. The 2nd derivative reveals this information. c"(t) = 24(t^2)(t^3 - 4)/(t^3 + 2)^3 Thus, t = 0, t = cuberoot(4) are critical points, making t = 1 a min or max. c"(1) < 0, making c(1) a max. NOW it's done.